OpenStudy (anonymous):

Find the difference of 9 and the square root of 7 and the square root of 28. A. 7 and the square root of 7 B. 11 and the square root of 7 C. 5 and the square root of 7

jimthompson5910 (jim_thompson5910):

\[\Large 9\sqrt{7}-\sqrt{28}\] \[\Large 9\sqrt{7}-\sqrt{4*7}\] \[\Large 9\sqrt{7}-\sqrt{4}*\sqrt{7}\] \[\Large 9\sqrt{7}-2\sqrt{7}\] \[\Large 7\sqrt{7}\]

OpenStudy (anonymous):

Thanks for helping me out!!

jimthompson5910 (jim_thompson5910):

np, glad to be of help

OpenStudy (anonymous):

(: Thanks soooo much (: you're awesome

jimthompson5910 (jim_thompson5910):

yw, anytime